OpenStudy (anonymous):

Help me step-by-step understand how to find the derivative of x^2

OpenStudy (anonymous):

f (x) = x^2 d (f (x)) / dx = d/dx (x^2)

OpenStudy (anonymous):

Therefore the answer equals 2x

OpenStudy (anonymous):

Can you go step by step, like how you got to an answer?

OpenStudy (anonymous):

I need to just watch videos and learn it lol

OpenStudy (anonymous):

I'm very sorry, it's hard to explain. Is it a question where you have to explain how you got the answer?

OpenStudy (jhannybean):

well, you could also understand the formula \(nx^{n-1}\) where n is the power of the variable you're derivitating.

OpenStudy (jhannybean):

So if you take n = 2 and follow the formula for derivatives, you'll get x^2 2 * x^{2-1} 2 * x^1 2x

OpenStudy (anonymous):

That's probably the easiest way to get the answer. ^

OpenStudy (anonymous):

Thanks, that helps a lot!